Ivanhoe Corporation wishes to exchange a machine used in its operations. Ivanhoe has received the following offers from other companies in the industry.

1. Shamrock Company offered to exchange a similar machine plus $25,760. (The exchange has commercial substance for both parties.)
2. Bridgeport Company offered to exchange a similar machine. (The exchange lacks commercial substance for both parties.)
3. Indigo Company offered to exchange a similar machine, but wanted $3,360 in addition to Ivanhoe’s machine. (The exchange has commercial substance for both parties.)


In addition, Ivanhoe contacted Sweet Corporation, a dealer in machines. To obtain a new machine, Ivanhoe must pay $104,160 in addition to trading in its old machine.

Ivanhoe

Shamrock

Bridgeport

Indigo

Sweet

Machine cost $179,200 $134,400 $170,240 $179,200 $145,600
Accumulated depreciation 67,200 50,400 79,520 84,000 –0–
Fair value 103,040 77,280 103,040 106,400 207,200


For each of the four independent situations, prepare the journal entries to record the exchange on the books of each company. (Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts.)

1.) Ivanhoe corp DR CR

-

-

-

-

-

Shamrock Corp DR CR

-

-

-

-

-

2.) Ivanhoe Corp DR CR

-

-

-

-

Brigeport Corp DR CR

-

-

-

3.) Ivanhoe Corp DR CR

-

-

-

-

-

Indigo Corp DR CR
-

-

-

-

-

4.) Ivanhoe Corp DR CR

-

-

-

-

-

Sweet Corp DR CR

-

-

-

To record exchange of inventory

-

-

To record cost of Inventory
Answer #1

1) Shamrock Company offered to exchange the similar Machine Machine Cost (a) Accumulated Depreciation (b) Book Value = (a-b)

2) Bridgeport Company offered to exchange the similar Machine Machine Cost (a) Accumulated Depreciation (b) Book Value = (a-b

3) Indigo Company offered to exchange the similar Machine Ivanhoe Machine Cost (a) Accumulated Depreciation (b) Book Value =

4) Sweet Company Sweet 145,600 Machine Cost (a) Accumulated Depreciation (b) Book Value = (a-b) Fair Value (d) Ivanhoe 179,20